OpenStudy (anonymous):

25^2c = square root of 5^(4c - 14)

OpenStudy (bohotness):

25^(2c) = (√5)^(4c - 14) ((√5)^4)^(2c) = (√5)^(4c - 14) (√5)^(8c) = (√5)^(4c - 14) 8c = 4c - 14 4c = -14 c = -14/4 = -7/2

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

@bohotness I have a question. How would you finish this problem if you went another route and did 25^4c = 5^(4c-14) for step two?

Nnesha (nnesha):

^

OpenStudy (bohotness):

sure

OpenStudy (bohotness):

mhmm lwet me think

OpenStudy (bohotness):

it would /2

Nnesha (nnesha):

answer is 2 ??